AIDA GELINA BRIKEN nToF CRIB ISOLDE CIRCE nTOFCapture DESPEC DTAS EDI_PSA 179Ta CARME StellarModelling DCF K40
  CARME  ELOG logo
Message ID: 127     Entry time: Mon Oct 18 16:25:11 2021
Author: JM RSS TD 
Subject: NEG rotation + cable removals 

Kapton cabling and the broken ceramic detector holder have been separated. Required holding jacking plate in a vice and using lots of propanol to wet the jacking screw. A long set of kapton cabling has been removed from the other side of the chamber which can be used when mounting the detectors. The feedthrough flange was re-mounted on that side with no kapton cabling. 

The wiring for the NEG feedthroughs was in confilct with detectors in previous position. Feedthroughs have been moved to the DN40 flanges at the top and bottom of the chamber section. In order to wire the NEGs to the feedthroughs required rotating the NEGs one position anti-clockwise and re-wiring some of the connections. 

Left two Getters are connected in series and are connected to the top feedthrough. The two right and the bottom getter are connected in series and are connected to the bottom feedthrough

Attachment 1: Screenshot_20211018-174141_Signal.jpg  166 kB  Uploaded Mon Oct 18 17:42:35 2021  | Hide | Hide all
Screenshot_20211018-174141_Signal.jpg
Attachment 2: 20211018_111130.jpg  1.360 MB  Uploaded Mon Oct 18 17:42:54 2021  | Hide | Hide all
20211018_111130.jpg
Attachment 3: 20211018_100015.jpg  1.130 MB  Uploaded Mon Oct 18 17:43:13 2021  | Hide | Hide all
20211018_100015.jpg
ELOG V3.1.4-unknown